OpenStudy (anonymous):

A standard deck of cards has had all the face cards removed so that only the ace through 10 of each suit remains. A card is drawn from the deck and a 6-sided die is rolled. Assume the Ace has a value of 1. Determine the probability of: a) drawing an even card (2, 4, 6, 8, or 10 of any suit) followed by an even die roll b) drawing an Ace following by die roll of 5 or 6

OpenStudy (anonymous):

For a) I got 1/2 for Even Card * 1/2 for Dice = 1/4 b) 4/40 for Ace * 2/6 for Dice = 1/30

OpenStudy (anonymous):

that is correct